# UNI: Unlearning-based Neural Interpretations

### v1

@@ 5.2 Robustness
Next, we evaluate UNI’s robustness to fragility adversarial attacks on model interpretations. Following Ghorbani et al. (2019a), we design norm-bounded attacks to maximise the disagreement in attributions whilst constraining that the prediction label remains unchanged. We consider a standard $l_{\infty}$ attack designed with FGSM (Goodfellow et al., 2014), with perturbation budget $\epsilon_{f}=8/255$.
$\delta_{f}^{*}=$ $\textrm{arg}\max_{\|\delta_{f}\|_{p}\leq\epsilon_{f}}\frac{1}{d_{X}}\sum_{i=1}^{d_{X}}d\left(\mathcal{A}(i,F,c,x),\mathcal{A}(i,F,c,x+\delta_{f})\right)$ (3)
@@
We report robustness results using 2 distance measures—Spearman correlation coefficient in Figure 4 and top-$k$ pixel intersection score in Figure 4—pre and post attack. While other methods like DeepLIFT (DL), BlurIG, Integrated Gradients (IG) are misled to output irrelevant feature saliencies, UNI robustly maintains attribution consistency and achieves the lowest attack attribution disagreement scores (before and after FGSM attacks) for both metrics.
@@
5.3 Stability
We compare UNI and other methods’ sensitivity to Riemann approximation noise, which manifests in visual artifacts and misattribution of salient features. As seen from Figures 6, 7, UNI reliably finds unlearned, “featureless" baselines for consistent attribution, regardless of the number of approximation steps $B\in\{1,15,30\}$. This is due to the low geodesic curvature of $\gamma^{\texttt{UNI}}$, which approximately minimises the local distance between points used in Riemann approximation.
@@ 6 Related Work
Perturbative methods perturb inputs to change and explain outputs (Sculley et al., 2015), including LIME (Ribeiro et al., 2016), SHAP, KernelSHAP and GradientSHAP (Lundberg et al., ), RKHS-SHAP (Chau et al., 2022), ConceptSHAP (Yeh et al., 2020), InterSHAP (Janzing et al., 2020), and DiCE (Kommiya Mothilal et al., 2021). LIME variants optimise a simulator of minimal functional complexity able to match the black-box model’s local behaviour for a given input-label pair. SHAP (Lundberg et al., ) consolidates LIME, DeepLIFT (Shrikumar et al., 2016), Layerwise Relevance Propagation (LRP) (Montavon et al., 2019) under the general, game-theoretic framework of additive feature attribution methods. For this framework, they outline the desired properties of local accuracy, missingness, consistency; they propose SHAP values as a feature importance measure which satisfies these properties under mild assumptions to generate model-agnostic explanations. However, such methods fail to give a global insight of the model’s decision function and are highly unstable due to the reliance on local perturbations (Fel et al., 2022). Bordt et al. (2022) show that this leads to variability, inconsistency and unreliability in generated explanations, where different methods give incongruent explanations which cannot be acted on.
@@
This family of post-hoc attributions is attractive due to its grounding in cooperative game-theory (Friedman, 2004). It comprises Integrated Gradients (Sundararajan et al., 2017), Adversarial Gradient Integration (Pan et al., 2021), Expected Gradients (EG) (Erion et al., 2021), Guided Integrated Gradients (GIG) (Kapishnikov et al., 2021) and BlurIG (Xu et al., 2020). Path attribution typically relies on a baseline – a “vanilla" image devoid of features; a path—an often linear path from the featureless baseline to the target image—along which the path integral is computed for every pixel. Granular control over the attribution process comes with difficulties of defining an unambiguously featureless baseline (for each (model, image) pair) (Sturmfels et al., 2020) and then defining a reliable path of increasing label confidence without intermediate inflection points (Akhtar & Jalwana, 2023). To measure the discriminativeness of features identified by attribution methods, experimental benchmarks such as ROAR (Hooker et al., 2019), DiffRAOR (Shah et al., 2021), RISE (Petsiuk et al., 2018) and the Pointing Game (Zhang et al., 2018) have been proposed.
@@ 7 Conclusion
In this work, we formally discuss the limitations of current path-attribution frameworks, outline a new principle for optimising baseline and path features, as well as introduce the UNI algorithm for unlearning-based neural interpretations. We empirically show that present reliance on static baselines imposes undesirable post-hoc biases which are alien to the model’s decision function. We account for and mitigate various infidelity, inconsistency and instability issues in path-attribution by defining principled baselines and conformant path features. UNI leverages insights from unlearning to eliminate task-salient features and mimic baseline activations in the “absence of signal". It discovers low-curvature, stable paths with monotonically increasing output confidence, which preserves the completeness axiom necessary for path attribution. We visually, numerically and formally establish the utility of UNI as a means to compute robust, meaningful and debiased image attributions. Future extensions to UNI include going beyond first-order approximate unlearning towards adopting certified, second-order machine learning techniques; as well as granular investigations on how the baseline definition and model’s robustness and inductive biases exert influence on path attribution results.
@@ Appendix A Appendix
We supplement the main text with visualisations of the UNI baseline, attributions and path features (properties, stability and robustness). We additionally include figures elucidating the colour, texture and frequency biases post-hoc imposed by path attribution methods. From Figure 7, we observe the stability of UNI path features: our attributions can be reliably and efficiently computed with Riemann approximation. In Figures 8, 9 and 10, we present visualisations on ImageNet-C, highlighting how static choices of baselines may bias the path-attribution procedure, leading to null or noisy explanations. UNI does not impose additional post-hoc assumptions that are alien to the model’s decision function. Furthermore, we present qualitative comparisons of attribution results of pre-trained models on the ImageNet-1K test set, in Figures 11, 12, 13, 14, 15 and 16. UNI attributions are visibly better localised and more semantically meaningful. Finally, we visualise the consistent, geodesic paths of monotonically increasing output confidence, discovered by UNI. As seen from Figures 17, 18, 19, 20, 21 and 22, while other path attribution methods might encounter extrema and turning points along the interpolation path from baseline to input, UNI’s path features are monotonic and preserve the crucial completeness property on which the path attribution framework depends.

### v2

@@ 5.2 Robustness
Next, we evaluate UNI’s robustness to fragility adversarial attacks on model interpretations. Following Ghorbani et al. (2019a), we design norm-bounded attacks to maximise the disagreement in attributions whilst constraining that the prediction label remains unchanged. We consider a standard $l_{\infty}$ attack designed with FGSM (Goodfellow et al., 2014), with perturbation budget $\varepsilon_{f}=8/255$.
$\delta_{f}^{*}=$ $\textrm{arg}\max_{\|\delta_{f}\|_{p}\leq\varepsilon_{f}}\frac{1}{d_{X}}\sum_{i=1}^{d_{X}}d\left(\mathcal{A}(i,F,c,x),\mathcal{A}(i,F,c,x+\delta_{f})\right)$ (3)
@@
We report robustness results using 2 distance measures—Spearman correlation coefficient in Table 6 and top-$k$ pixel intersection score in Table 6—pre and post attack. While other methods like DeepLIFT (DL), BlurIG, Integrated Gradients (IG) are misled to output irrelevant feature saliencies, UNI robustly maintains attribution consistency and achieves the lowest attack attribution disagreement scores (before and after FGSM attacks) for both metrics.
5.3 Stability
We compare UNI and other methods’ sensitivity to Riemann approximation noise, which manifests in visual artefacts and misattribution of salient features. As seen from Figures 7, 10, UNI reliably finds unlearned, “featureless" baselines for consistent attribution, regardless of the number of approximation steps $B\in\{1,15,30\}$. This is due to the low geodesic curvature of $\gamma^{\texttt{UNI}}$, which approximately minimises the local distance between points in Riemann approximation.
@@ 6 Related Work
Perturbative methods perturb inputs to change and explain outputs (Sculley et al., 2015), including LIME (Ribeiro et al., 2016), SHAP, KernelSHAP and GradientSHAP (Lundberg et al., ), RKHS-SHAP (Chau et al., 2022), ConceptSHAP (Yeh et al., 2020), InterSHAP (Janzing et al., 2020), and DiCE (Kommiya Mothilal et al., 2021). LIME variants optimise a simulator of minimal functional complexity able to match the black-box model’s local behaviour for a given input-label pair. SHAP (Lundberg et al., ) consolidates LIME, DeepLIFT (Shrikumar et al., 2016), Layerwise Relevance Propagation (LRP) (Montavon et al., 2019) under the general, game-theoretic framework of additive feature attribution methods. For this framework, they outline the desired properties of local accuracy, missingness, consistency; they propose SHAP values as a feature importance measure which satisfies these properties under mild assumptions to generate model-agnostic explanations. However, such methods fail to give a global insight of the model’s decision function and are highly unstable due to the reliance on local perturbations (Fel et al., 2022b). Bordt et al. (2022) show that this leads to variability, inconsistency and unreliability in generated explanations, where different methods give incongruent explanations which cannot be acted on. Recent works have made considerable progress, including RISE (Petsiuk et al., 2018), which strives to causally explain model predictions by approximating the necessary saliency of pixels through random masking; Sobol (Fel et al., 2021), which adapts Sobol indices for perturbation masks towards variance-based sensitivity analysis; and FORGrad (Muzellec et al., 2023), which filters out high-frequency gradient noise induced by white-box methods (and network pooling or striding operations) and which can be complementarily applied to further UNI’s explanation faithfulness and efficiency.
@@
This family of post-hoc attributions is attractive due to its grounding in cooperative game-theory (Friedman, 2004). It comprises Integrated Gradients (Sundararajan et al., 2017), Adversarial Gradient Integration (Pan et al., 2021), Expected Gradients (EG) (Erion et al., 2021), Guided Integrated Gradients (GIG) (Kapishnikov et al., 2021) and BlurIG (Xu et al., 2020). Path attribution typically relies on a baseline – a “vanilla" image devoid of features; a path—an often linear path from the featureless baseline to the target image—along which the path integral is computed for every pixel. Granular control over the attribution process comes with difficulties of defining an unambiguously featureless baseline (for each (model, image) pair) (Sturmfels et al., 2020) and then defining a reliable path of increasing label confidence without intermediate inflection points (Akhtar & Jalwana, 2023). To measure the discriminativeness of features identified by attribution methods and the extent to which model predictions depend on them, experimental benchmarks and metrics such as ROAR (Hooker et al., 2019), DiffRAOR (Shah et al., 2021), deletion/insertion score (Petsiuk et al., 2018), the Hilbert-Schmidt independence criterion (HSIC) (Novello et al., 2022) and the Pointing Game (Zhang et al., 2018a) have been proposed.
@@ 7 Conclusion
In this work, we formally discuss the limitations of current path-attribution frameworks, outline a new principle for optimising baseline and path features, as well as introduce the UNI algorithm for unlearning-based neural interpretations. We empirically show that present reliance on static baselines imposes undesirable post-hoc biases which are alien to the model’s decision function. We account for and mitigate various infidelity, inconsistency and instability issues in path-attribution by defining principled baselines and conformant path features. UNI leverages insights from unlearning to eliminate task-salient features and mimic baseline activations in the “absence of signal". It discovers low-curvature, stable paths with monotonically increasing output confidence, which preserves the completeness axiom necessary for path attribution. We visually, numerically and formally establish the utility of UNI as a means to compute robust, meaningful and debiased image attributions.
The contributions of UNI extend beyond the presented method and analyses, towards investigating machine unlearning as a tool for white-box interpretability. Unlearning at different granularities allows us to audit the various levels of a model’s learned feature hierarchy. In this work, we illustrate how first-order, sample-wise unlearning can identify salient input features important for a single prediction. A promising future direction involves interpreting higher-level, semantically complex concepts required to learn a task or fit a data distribution, by instead unlearning a set of concept-clustered exemplars. It is also of interest to delve into how interpretability methods impose additional assumptions onto trained models, prompting questions such as how to best design and align the correct interpretability method for a given model; how to use attribution methods to compare and contrast the inductive biases of different network architectures, of models trained with robust versus non-robust objectives, of models trained using different equivariant data augmentation strategies. Further technical extensions to UNI include going beyond first-order approximate unlearning towards certified, second-order machine unlearning; as well as granular investigations of how the baseline definition, model’s robustness and model’s inductive biases exert influence on path attribution results.
@@ Appendix A Appendix
A.1 Verifying Attribution Specificity
To verify that UNI computes explanations that are specific to each task–model–input triplet, we compare its saliency attributions across models for the same image input. Visually, we observe in Figure 8 that attributions differ significantly and even reflect the inductive biases of respective models (e.g. grid-like artefacts are present in ViT attributions whereas smoother attributions are computed for convolutional architectures). We further present numerical results in Table 8—LPIPS (Zhang et al., 2018b) scores reflect the dissimilarity/distance between the original image and the unlearned baseline; the percentage change in confidence scores reflect how the unlearned baseline effectively reduces predictive confidence (relative to the original input).
[…]
ResNet-18 $82.3\%$$021$$\pm.025$$003$$\pm.005$
Eff-v2-s $76.9\%$$025$$\pm.024$$004$$\pm.011$
ConvNeXt-T $95.1\%$$018$$\pm.016$$002$$\pm.003$
VGG-16-bn $71.6\%$$017$$\pm.020$$001$$\pm.002$
ViT-B-16 $69.7\%$$014$$\pm.015$$004$$\pm.007$
SwinT $84.6\%$$014$$\pm.015$$002$$\pm.002$
Figure 8: UNI computes different attributions to explain the predictions of each model.
A.2 Preliminary Results on NLP
Table 8: Faithfulness: $L_{2}$-Distance from activation patching to attribution patching results on the residual stream (averaged over 100 samples).
UNI Random
Pythia-1b-v0 $\bm{3.12}$ $6.64$
GPT2-medium $\bm{15.26}$ $35.00$
Llama-3.2-1B $\bm{5.25}$ $10.17$
Figure 9: Visual comparison of attribution results for Activation vs. Attribution patching, with UNI versus Random baselines, on Pythia-1b-v0. Each cell shows the logit variation obtained by patching at a specific token and layer the residual stream of our baseline with the original activation.
We extend the testing of our method to the case of Natural Language Processing (NLP). We choose to test the application of UNI in the general framework of generative models (which includes classification models), and attribution of not only inputs but more generally activations. Activation patching (Heimersheim & Nanda, 2024) is one of the most widely used technique in Mechanistic Interpretability, and more generally to study the properties of LLMs’ internals (Vig et al., 2020; Meng et al., 2023; Wang et al., 2022; Feng & Steinhardt, 2024; Cunningham et al., 2023; Stolfo et al., 2023; Hanna et al., 2023). This attribution method consists in analyzing a model’s output variation after replacing its internal activations, following the equation:
$\mathcal{A}_{e}^{\textsc{Act}}(x)=F(x|e=e(x^{\prime}))-F(x)$ (4)
where $e$ denotes one activation in the model, $x^{\prime}$ a chosen baseline, and $F$ a function of the model’s output (usually the logit value of the maximum probability token of the model run on $x$).
Unfortunately activation patching is computationally costly, especially for purposes such as circuit discovery (Conmy et al., 2023). One of the main alternatives that solves the scalability problem is attribution patching (Nanda, 2023; Syed et al., 2023), which computes a first order Taylor approximation of Equation 4:
$\mathcal{A}_{e}^{\textsc{Attr}}(x)=(e(x)-e(x^{\prime}))^{T}\nabla_{e}F(x^{\prime})$ (5)
for which the attributions for all of the activations can be computed at the same time (no patching of one single activation is performed). Despite its scalability, attribution patching suffers from a lack of faithfulness for causal interventions, mainly due to saturation and lack of linearity of the studied dependencies.
The analogy with integrated gradients seems quite striking, and indeed two recent works (to our knowledge) have tried to investigate the use of IG for more faithful attribution patching. While Marks et al. (2024) applies a very computationally complex version of such a method to small models, Hanna et al. (2024) proves the potential of IG-based attribution patching, while showing it still gets outperformed by activation patching.
We here provide a new UNI-based attribution method algorithm outputting faithful attributions while maintaining the scalability advantage of attribution patching. Mainly, we apply Algorithm 1 to compute a baseline $x^{\prime}$ that is then used to compute Equation 1:
[…]
where we take $x$ to be the embedding of the input, to allow for continuous operations on it. Considering the known high faithfulness of activation patching (Hanna et al., 2024), we approximate faithuflness of attributions computed from a baseline, as the $L_{2}$-distance between these attributions and the activation patching ones. The dataset used is a subset of $100$ counterfactual prompts taken from Meng et al. (2023), and three different models are tested: Pythia-1b-v0 (Biderman et al., 2023), GPT2-medium (Radford et al., 2019) and Llama-3.2-1B (Dubey et al., 2024). The results can be seen in Table 8, and visuals in Figure 9. Note that no fine-tuning of UNI hyperparameters has been done, so that we expect even better results when adapting for each models. Unsurprisingly, decoding the baselines by shortest distance to the rows of the embedding matrix yields the same input, and a direct decoding of the perturbation $\delta$ doesn’t provide any interesting information.
A.3 Additional Visualisations
@@
We supplement the main text with visualisations of the UNI baseline, attributions and path features (properties, stability and robustness). We additionally include figures elucidating the colour, texture and frequency biases post-hoc imposed by path attribution methods. From Figure 10, we observe the stability of UNI path features: our attributions can be reliably and efficiently computed with Riemann approximation. In Figures 11, 12 and 13, we present visualisations on ImageNet-C, highlighting how static choices of baselines may bias the path-attribution procedure, leading to null or noisy explanations. UNI does not impose additional post-hoc assumptions that are alien to the model’s decision function. Furthermore, we present qualitative comparisons of attribution results of pre-trained models on the ImageNet-1K test set, in Figures 14, 15, 16, 17, 18 and 19. UNI attributions are visibly better localised and more semantically meaningful. Finally, we visualise the consistent, geodesic paths of monotonically increasing output confidence, discovered by UNI. As seen from Figures 20, 21, 22, 23, 24 and 25, while other path attribution methods might encounter extrema and turning points along the interpolation path from baseline to input, UNI’s path features are monotonic and preserve the crucial completeness property on which the path attribution framework depends.